jennylove:

An electric field is measured at 3.00×105 N/C at a distance from a source point charge of 3.00×10−6 C. How far is the point from the source charge? 0.09 m 0.30 m 0.36 m 0.60 m

jennylove:

im not sure where to begin :/

justjm:

You must use the formula \(E=k\frac{Q}{r^2}\) where k=Coulomb's constant E=Electric field in N/C Q=Point charge in C r=radius in m Plugging in the values \(3.00\times10^5~\text{N/C}=k\frac{3.00\times10^{-6}}{r^2}\) You can use algebra to isolate r

justjm:

Did you understand it yet? for 'k' you might need to get that online or on your calculator